Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 426     Entry time: 30 Oct 2023, 23:18
Author: Jan Gruenert 
Group: PRC 
Subject: Status 

At the end of the MON late shift, the tuning status and obtained lasing pulse energies (ATT#1) are as follows:

  • SA1 at 19 keV : 660 uJ +/- 23 uJ
  • SA2 at 18 keV : 1073 uJ +/- 55 uJ
  • SA3 at 1.2 keV : 4330 uJ +/- 122 uJ

SD for a stable 2 min interval, fast XGM train average (INTENSITY.RAW.TRAIN)

More tuning will happen overnight, and for SA3 the photon energies 2.5keV and 3.0 keV will also be prepared.

Attachment 1: JG_2023-10-30_um_23.15.36.png  834 kB  | Hide | Hide all
JG_2023-10-30_um_23.15.36.png
ELOG V3.1.4-7c3fd00